OpenStudy (anonymous):

In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rat b. Use the table to determine at approximately which point the number of cells will be the same for each organism. c. Graph the system of equations, and show the point of intersection. d.Explain what the points graphed for each line represent. e.Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph from part c. f.Find the point(s) of intersection, and explain what the intersection represents in the context of the problem.

OpenStudy (anonymous):

S(X)=100+23X M(X)= 90(1.2*)

OpenStudy (anonymous):

@wio